Find the sum of: (a) 137 and – 354 (b) – 52 and 52  
owerbhierhitom,niomnrimmoemhnteioanmeioam,irohomtei

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(a) 137 and -354 (137) + (-354)
= (317) + (-137) + (-217)
= 0 + (-217) [(317) + (-317)= 0]
= (-217)

(b) -52 and 52
(-52) + (+52) = 0
[(-a) + (+a) = 0]


Thanks